A billiard ball moving at 6.00 m/s strikes a stationary ball of the same mass. After the collision, the first ball moves at 5.20 m/s, at an angle of +30.0
Answer #1
given:

moving ball:
mass = m
vi = 6 m/s

stationary ball:
mass = m
vi = 0 m/s

conservation of momentum:
x-direction:
(m)(6) + (m)(0) = m(5.2cos30) + m(vfx)
6m = 4.5m + m(vfx)
vfx = 1.50 m/s

y-direction:
m(0) + m(0) = m(5.2sin30) + m(vfy)
-2.6m = mvfy
vfy = -2.6 m/s

magnitude:
vf = sqrt[vfx^2 + vfy^2]
vf = 3.00 m/s

direction:
arctan(vfy/vfx) = -60.02 degrees

BOL